OpenStudy (anonymous):

two terms of an arithmetic sequence are t5=-31.4 and t10=-50.4. what is t23?

OpenStudy (anonymous):

t10-t5 equal -19, 10-5=5 so t, -19/5=-3.8, is t1 so t32 is -3.8times 32

OpenStudy (anonymous):

when I multiply -3.8 to 32, I come up with -121.6 but -121.6 isn't one of the given choices

OpenStudy (anonymous):

should be t10+t5

OpenStudy (campbell_st):

set up 2 equations using the formula for a term in an AP \[a_{n} = a + (n - 1)d\] a = 1st term, n = number of terms, d = common difference so solve simultaneously term 10: a + 9d = -50.4 equation 1 term 5 a + 4d = -31.4 equation 2 --------------- eq 1 - 2 5d = -19 so the common difference is d = -3.8 substitute into either equation to find the 1st term when you get that you can use the formula above to find the 23 term, n = 23

OpenStudy (amistre64):

an arithmetic sequence is a discrete form of a line .... you are given 2 points to define the line with